# Faster and Better LLMs via Latency-Aware Test-Time Scaling

### v3

@@ Abstract
Zili Wang1,2thanks: Equal contribution., Tianyu Zhang3∗, Lei Zhu3,†, Haoli Bai3, Lu Hou3, Shiming Xiang1,2,†, Xianzhi Yu3, Wulong Liu3 1 School of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing, China 2 Institute of Automation, Chinese Academy of Sciences, China 3 Huawei Noah’s Ark Lab wangzili2022@ia.ac.cn, {zhangtianyu59,zhulei168}@huawei.com
@@ 1 Introduction
With the number of generated tokens (#tokens) as budget, many existing studies Snell et al. (2024); Setlur et al. (2025); Yang et al. (2025); Liu et al. (2025); Shi and Jin (2025); Zhang et al. (2024) have examined compute-optimal strategies that enhance average performance gain per token, a metric we refer to as token efficiency. However, in latency-sensitive scenarios where small batch sizes are employed, “compute-optimal” does not necessarily translate to “latency-optimal”. This discrepancy is because the performance achieved within a limited time is determined by token efficiency as well as the throughput (i.e., average number of output tokens per second). As shown in Figure 1, for s1.1-32B Muennighoff et al. (2025) model, although sequential scaling has a better token efficiency, achieving a similar performance with 9$\times$ fewer tokens compared to parallel scaling, it turns out that parallel scaling achieves 1.6$\times$ lower latency. In fact, under small batch sizes, the time of an LLM autoregressive decoding step is dominated by the memory access of the parameters. Therefore, a moderately increased number of parallel branches incurs little additional latency, allowing a much higher throughput almost for free, as shown in Figure 3.
@@
To bridge the gap, we examine the impact of concurrency configuration for latency-aware test-time scaling in this study. Specifically, we conduct experiments on representative datasets including MATH-500 Hendrycks et al. (2021), AIME24 AoPS (2024), AIME25 AoPS (2025), and GPQA-Diamond Rein et al. (2024) across model sizes from 3B to 32B and under varied concurrency configurations. Revealed by our experiments and analyses, the latency-optimal concurrency depends on the comparative advantage of token efficiency in sequential scaling and parallel scaling. For models with higher token efficiency with parallel scaling (e.g., LLaMa-3.1-8B-Instruct Meta (2024)), one should prioritize branch-wise parallelism. Otherwise (e.g., for QwQ-32B Team (2025)) sequence-wise parallelism takes priority. Besides, under a latency constraint of 1 minute, an optimal concurrency configuration can provide up to 7.3% better performance than baselines where only a single parallelism is applied.
@@
    We provide a unified view for parallel branches and speculative decoding from the perspective of generation concurrency. This allows us to frame latency-optimal test-time scaling as a resource allocation problem.
@@
    Through extensive experiments, we explore the optimal concurrency configuration for latency-aware test-time scaling. Our experiments reveal that for s1.1-32B, an optimal concurrency configuration can improve the accuracy by 7.3% while reducing latency by 1.7$\times$, reaching 82.3% on MATH-500 in 1 minute.
@@ 2 Related Work
Scaling the compute at inference time has been proven to be a prominent approach to improve the performance of LLMs. Generally, test-time scaling methods fall into two main categories: sequential and parallel scaling. For the former, sequential scaling methods, represented by OpenAI’s o1 OpenAI (September 2024), DeepSeek-R1 Guo et al. (2025), and Qwen QwQ Team (2025), enforce the model to generate longer solutions with a detailed reasoning chain. Such an ability can be incentivized with supervised training Nye et al. (2021); Lee et al. (2025); Muennighoff et al. (2025) or reinforcement learning Guo et al. (2025). Another main category is parallel scaling. This method first generates multiple response candidates in parallel, then applies a selection criterion to identify the best output. Recent research primarily uses token count as a metric to measure the budget of test-time scaling Snell et al. (2024); Setlur et al. (2025); Yang et al. (2025). Different from previous works, our work points out that compute-optimal does not necessarily translate to latency-optimal. We introduce latency-aware test-time scaling and figure out the concurrency configuration to latency-optimal TTS.
@@
Speculative decoding uses a small draft model to generate several draft tokens autoregressively, and the target model to verify them in parallel while ensuring a lossless acceleration Leviathan et al. (2023); Chen et al. (2023). Many studies have proposed improvements to speculative decoding. EAGLE-series Li et al. (2024b, a, 2025), Medusa Cai et al. (2024), and Hydra Ankner et al. (2024) employ features from the target model for better draft acceptance rate. HASS Zhang et al. (2025) mitigates the inconsistency between training and inference by simulating the multi-step draft generation in the training phase. Our work demonstrates the feasibility of latency-aware TTS optimization through speculative decoding, presenting it as one viable approach to latency-optimal TTS.
@@ 3.1 Latency-Aware Test-time Scaling
Consider s1.1-32B Muennighoff et al. (2025) as an example. Figure 1 compares the accuracy-#token and accuracy-latency curves of sequential and parallel scaling. Sequential scaling achieves higher token efficiency (83.3%/$2^{11.3}$toks) than parallel scaling (83.3%/$2^{14.4}$toks), but Figure 3 reveals its throughput is 16×16$\times$16 × lower. Consequently, sequential scaling requires 1.6×1.6$\times$1.6 × more time to attain comparable accuracy. This suggests that while sequential scaling improves token efficiency, its inferior throughput hinders better accuracy in limited time. Thus, test-time scaling budgets must consider both token efficiency and throughput to achieve high accuracy with low latency.
@@ 4.2 Sequential and Parallel Scaling under Latency-Aware TTS
As shown in Figure 2, merely employing parallel scaling can surpass sequential scaling within a shorter, limited time. However, the extent of performance improvement varies depending on the model type. Specifically, for reasoning models like QwQ-32B, parallel scaling reaches $80.1\%$ accuracy on the MATH-500 dataset within $30$ seconds, but sequential scaling requires 1.4×1.4$\times$1.4 × more time to achieve the comparable performance. Sequential scaling shows a slight performance gain for LLaMa-3.1-8B-Instruct, which is not designed for long CoT reasoning. While parallel scaling is still effective on MATH-500 since it explores more diverse solution paths, it shows obvious token efficiency, and sequential scaling cannot surpass it. Overall, parallel scaling can surpass sequential scaling w.r.t. latency, reaching a comparable accuracy within a relatively short time.
@@ 4.4 The Impact of Sequence-wise Parallelism for Latency-Aware TTS
Speculative decoding can accelerate model inference to some extent while preserving accuracy. As shown in Figure 6, most scaling curves with speculative decoding are on the left side of the baseline curve. As the speed-up ratio grows, the left-shifted trend becomes more obvious. However, when the draft length $\gamma$ becomes large, the speed-up ratio would decay, slowing down the overall speed. In our experiments, QwQ-32B is the target model, while DeepSeek-R1-Distill-Qwen-7B as the draft model, achieving a maximum speed-up ratio of 1.64×1.64$\times$1.64 ×. Therefore, the scaling curve moves to the left first and then to the right. Sequential scaling with speculative decoding can achieve $7.5\%$ higher accuracy than the baseline curve for a limited time. These results demonstrate that employing a proper draft length $\gamma$ can push the inference speed to an optimal stage, enabling further sequential scaling within a limited time.
@@ 5.2 How Can Latency-Optimal Improve Compared with Baseline?
Based on previous findings of latency-optimal TTS with different model types, branches, and draft lengths, we summarize the results in Table 1. For s1.1-32B on MATH-500, we find that latency-optimal TTS can achieve $6\%$ accuracy improvement on average than merely using sequence-wise parallelism, and 1.6×1.6$\times$1.6 × on average faster than branch-wise parallelism. Latency-optimal TTS outperforms the baseline in both accuracy and latency to a noticeable degree. However, increased branch and draft length causes LLMs to enter a compute-bound regime, where computational overhead exceeds memory access overhead. This eliminates the benefits of parallelism, leading to increased latency and sometimes even worse than the baseline. This suggests that a latency-optimal TTS strategy requires extremely fine-grained parameter tuning.
@@ 6 Conclusion
In this paper, we propose to rethink test-time scaling in latency-sensitive scenarios. We show that compute-optimal does not always result in latency-optimal under such conditions due to memory-bound constraint. To address this, we propose to improve TTS on generation concurrency to maximize throughput from a unified view. Specifically, we present two approaches: (1) branch-wise parallelism via multiple branches and (2) sequence-wise parallelism by speculative decoding, along with their combinations. Furthermore, we investigate the concurrency allocation strategy to balance these approaches for latency-optimal TTS. Experimental results show that latency-optimal TTS enables 32B model to achieve 82.3% accuracy on MATH-500 within 1 minute, while 3B model attains notable accuracy within just 10 seconds, showing significant improvements in both speed and accuracy.
@@ 7 Limitations
The maximum sequence length in our experiment settings is 8192. Note that, as the number of branches and sequence length increase, the substantial memory access burden imposed by the KV cache leads to a significant rise in latency. This causes methods that increase concurrency to experience corresponding increases in latency. Additionally, longer sequence lengths may introduce performance degradation due to excessive context length in the model Muennighoff et al. (2025). Since our work adopts relatively common experimental configurations, these issues have not manifested prominently, and thus our proposed method and conclusions remain valid.
@@ A.2 Additional Results
We present additional experimental results. The results on the influence of branch-wise parallelism is shown in Figure 11. The results of the influence of branch-wise parallelism under different speculative decoding configurations is shown in Figure 12. The results of sequence-wise parallelism under different branch counts are shown in Figure 13. The results of the combined branch-wise parallelism and sequence-wise parallelism configurations w.r.t. fixed generation concurrency $F$ are shown in Figure 14.
@@
Figure 14: Results of different branch-wise parallelism and sequence-wise parallelism configurations with fixed generation concurrency.

### v4

@@ Abstract
Zili Wang1,2thanks: Equal contribution., Tianyu Zhang3∗, Haoli Bai3, Lu Hou3, Xianzhi Yu3, Wulong Liu3, Shiming Xiang1,2,†, Lei Zhu3,† 1 School of Artificial Intelligence, University of Chinese Academy of Sciences, China 2 MAIS, Institute of Automation, Chinese Academy of Sciences, China 3 Huawei Noah’s Ark Lab wangzili2022@ia.ac.cn, {zhangtianyu59,zhulei168}@huawei.com
@@ 1 Introduction
[…]
@@
To bridge the gap, we examine the impact of concurrency configuration for latency-aware test-time scaling in this study. Specifically, we conduct experiments on representative datasets including MATH-500 Hendrycks et al. (2021), AIME24 AoPS (2024), AIME25 AoPS (2025), and GPQA-Diamond Rein et al. (2024) across model sizes from 3B to 32B and under varied concurrency configurations. Revealed by our experiments and analyses, the latency-optimal concurrency depends on the comparative advantage of token efficiency in sequential scaling and parallel scaling. For models with higher token efficiency with parallel scaling (e.g., LLaMa-3.1-8B-Instruct Meta (2024)), one should prioritize branch-wise parallelism. Otherwise (e.g., for QwQ-32B Team (2025)) sequence-wise parallelism takes priority. To determine the latency-optimal configuration, we propose a simple yet effective greedy search algorithm with less searching steps. Under 1 minute latency constraint, an optimal concurrency configuration can provide up to 7.3% better performance than baselines where only a single parallelism is applied.
@@
    We provide a unified view for parallel branches and speculative decoding from the perspective of generation concurrency. This allows us to frame latency-optimal test-time scaling as a resource allocation problem. A greedy search algorithm is proposed to search the latency-optimal configuration.
@@
[…]
@@ 2 Related Work
Scaling the compute at inference time has been proven to be a prominent approach to improve LLM performance. Generally, test-time scaling methods fall into two main categories: sequential and parallel scaling. For the former, sequential scaling methods, represented by OpenAI’s o1 OpenAI (September 2024), DeepSeek-R1 Guo et al. (2025), and Qwen QwQ Team (2025), enforce the model to generate longer solutions with a detailed reasoning chain. Such an ability can be incentivized with supervised training Nye et al. (2021); Lee et al. (2025); Muennighoff et al. (2025) or reinforcement learning Guo et al. (2025). Another main category is parallel scaling. This method first generates multiple response candidates in parallel, then applies a selection criterion to identify the best output. Recent research primarily uses token count as a metric to measure the budget of test-time scaling Snell et al. (2024); Setlur et al. (2025); Yang et al. (2025). Recent work Singhi et al. (2025) shows that majority voting outperforms verifier-based methods under low/medium token budgets. Notably, low latency necessitates limited token usage, where majority voting is superior. Thus, our work employs majority voting for parallel scaling. Different from previous works, our work points out that compute-optimal does not necessarily translate to latency-optimal. We introduce latency-aware test-time scaling and figure out the concurrency configuration to latency-optimal TTS.
@@
Speculative decoding uses a small draft model to generate several draft tokens autoregressively, and the target model to verify them in parallel while ensuring a lossless acceleration Leviathan et al. (2023); Chen et al. (2023). Many studies focus on improving acceleration. EAGLE-series Li et al. (2024b, a, 2025), Medusa Cai et al. (2024), and Hydra Ankner et al. (2024) employ features from the target model for better draft acceptance rate. HASS Zhang et al. (2025) mitigates the inconsistency between training and inference by simulating the multi-step draft generation in the training phase. Hierarchical Drafting Cho et al. (2025) improves the acceptance rate by regarding the hierarchical drafting strategy based on temporal locality. MoESD Huang et al. (2025) shows that at medium batch sizes, MoE models with speculative decoding can surpass dense models. Our work demonstrates the feasibility of latency-aware TTS optimization through speculative decoding, presenting it as one viable approach to latency-optimal TTS.
@@ 3.1 Latency-Aware Test-time Scaling
[…]
@@ 3.3 Latency-Optimal Test-Time Scaling
Algorithm 1 Greedy Search for Latency-Optimal TTS
1:tts_tasktts\_task: TTS evaluation. $T$: latency budget. $(B,\gamma)$: configuration; accacc: accuracy
[…]
6:while not endend do
7:  $acc_{b}\leftarrow tts\_task(2B,\gamma,T)$ $\triangleright$ Expand branch-wise
8:  $acc_{\gamma}\leftarrow tts\_task(B,\gamma+1,T)$ $\triangleright$ Expand sequence-wise
9:  if $acc_{b}\geq acc_{\gamma}$ then
10:   $B_{new},\gamma_{new},acc_{new}\leftarrow 2B,\gamma,acc_{b}$
11:  else
12:   $B_{new},\gamma_{new},acc_{new}\leftarrow B,\gamma+1,acc_{\gamma}$
13:  end if
14:  if $acc_{new}>acc$ then
15:   $B,\gamma,acc\leftarrow B_{new},\gamma_{new},acc_{new}$
16:  else
17:   $end\leftarrow\textbf{True}$
18:  end if
19:end while
20:return $(B,\gamma),acc$
Finding $(B^{*},\gamma^{*})$ requires to search the configuration space. Although grid search can find the optimal configuration, it brings a significant cost to search the entire configuration space. Therefore, we introduce a simple yet effective greedy search algorithm, as shown by Algorithm 1.
In practice, $B$ takes values in powers of 2 ($B=2^{k}$), while $\gamma$ increments in steps of 1. The basic grid search evaluates every $(B,\gamma)$, with complexity of $O((\log B_{max}+1)\times\gamma_{max})$. With greedy search, the complexity reduces to $O((\log B^{*}+1)+\gamma^{*})$, where $B^{*}$ and $\gamma^{*}$ denote the optimal configuration.
@@ 4.2 Sequential and Parallel Scaling under Latency-Aware TTS
As shown in Figure 2, merely employing parallel scaling can surpass sequential scaling within a shorter, limited time. However, the extent of performance improvement varies depending on the model type. Specifically, for reasoning models like QwQ-32B, parallel scaling reaches $80.1\%$ accuracy on the MATH-500 dataset within $30$ seconds, but sequential scaling requires $1.4\times$ more time to achieve the comparable performance. Sequential scaling shows a slight performance gain for LLaMa-3.1-8B-Instruct, which is not designed for long CoT reasoning. While parallel scaling is still effective on MATH-500 since it explores more diverse solution paths, it shows obvious token efficiency, and sequential scaling cannot surpass it. Overall, parallel scaling can surpass sequential scaling w.r.t. latency, reaching a comparable accuracy within a relatively short time.
@@ 4.4 The Impact of Sequence-wise Parallelism for Latency-Aware TTS
Speculative decoding can accelerate model inference to some extent while preserving accuracy. As shown in Figure 6, most scaling curves with speculative decoding are on the left side of the baseline curve. As the speed-up ratio grows, the left-shifted trend becomes more obvious. However, when the draft length $\gamma$ becomes large, the speed-up ratio would decay, slowing down the overall speed. In our experiments, QwQ-32B is the target model, while DeepSeek-R1-Distill-Qwen-7B as the draft model, achieving a maximum speed-up ratio of $1.64\times$. Therefore, the scaling curve moves to the left first and then to the right. Sequential scaling with speculative decoding can achieve $7.5\%$ higher accuracy than the baseline curve for a limited time. These results demonstrate that employing a proper draft length $\gamma$ can push the inference speed to an optimal stage, enabling further sequential scaling within a limited time.
@@ 4.5 Latency-Optimal TTS with Branch-wise and Sequence-wise Parallelism
Model Grid Search Greedy Search
$(B^{*},\gamma^{*})$ Accuracy Steps $(B^{*},\gamma^{*})$ Accuracy Steps
s1.1-32B $(16,5)$ 82.3% 56 $(16,5)$ 82.3% 10
s1.1-3B $(32,5)$ 72.4% 56 $(32,5)$ 72.4% 11
DS-32B $(16,4)$ 62.2% 56 $(16,4)$ 62.2% 9
QwQ-32B $(4,5)$ 94.5% 56 $(4,5)$ 94.5% 8
Table 2: Comparison of Grid Search and Greedy Search for Different Models. DS-32B: DeepSeek-Distill-32B.
@@ 5.2 How Can Latency-Optimal Improve Compared with Baseline?
Based on previous findings of latency-optimal TTS with different model types, branches, and draft lengths, we summarize the results in Table 1. For s1.1-32B on MATH-500, we find that latency-optimal TTS can achieve $6\%$ accuracy improvement on average than merely using sequence-wise parallelism, and $1.6\times$ on average faster than branch-wise parallelism. Latency-optimal TTS outperforms the baseline in both accuracy and latency to a noticeable degree. However, increased branch and draft length causes LLMs to enter a compute-bound regime, where computational overhead exceeds memory access overhead. This eliminates the benefits of parallelism, leading to increased latency and sometimes even worse than the baseline. This suggests that a latency-optimal TTS strategy requires extremely fine-grained parameter tuning.
5.3 Effectiveness of Greedy Search
To validate the effectiveness of our proposed greedy search algorithm, extensive experiments are conducted across various models on MATH-500. The results in Table 2 show that our greedy search achieves the same optimal configuration as grid search, while significantly reducing searching steps (8–10 vs 56). These results show the practical efficiency of our greedy search algorithm.
Aggregation Method 1024 2048 4096 8192
Acc. (%) Acc. (%) Acc. (%) Acc. (%)
Majority Voting 82.3 $\pm$1.8 85.9 $\pm$0.7 91.7 $\pm$1.3 92.7 $\pm$0.3
Confidence min-max 79.2
$\pm$1.6
82.2
$\pm$1.5
88.6
$\pm$1.4
90.1
$\pm$1.3
Confidence avg-max 81.2
$\pm$1.4
83.2
$\pm$1.4
90.6
$\pm$0.3
91.1
$\pm$1.2
Confidence min-vote 82.4
$\pm$1.3
85.2
$\pm$1.5
91.5
$\pm$1.6
92.7
$\pm$0.4
Confidence avg-vote 83.0 $\pm$1.2 86.3 $\pm$1.3 92.3 $\pm$1.2 93.5 $\pm$0.8
Table 3: Performance comparison of majority voting and different confidence-based strategies. Acc.: Accuracy. Results are obtained from 3 repeated experiments with mean and standard deviation reported.
5.4 Different Aggregation Strategies
We conduct additional ablation experiments with s1.1-32B on MATH-500 comparing majority voting (self-consistency) with confidence-based aggregation strategies Liu et al. (2025): first weight the answers by minimum (min) or average (avg) confidence, then select the answer with highest confidence (max) or accumulates the scores of all identical answers and then selects the answer with the highest score (vote). Results are shown in Table 3. Note that confidence-based strategies do not affect the latency. The results show that using confidence scores can indeed slightly improve output quality over simple majority voting.
@@ 6 Conclusion
In this paper, we propose to rethink test-time scaling in latency-sensitive scenarios. We show that compute-optimal does not always result in latency-optimal under such conditions due to memory-bound constraint. To address this, we propose to improve TTS on generation concurrency to maximize throughput from a unified view. Specifically, we present two approaches: (1) branch-wise parallelism via multiple branches and (2) sequence-wise parallelism by speculative decoding, along with their combinations. Furthermore, we investigate the concurrency allocation strategy to balance these approaches for latency-optimal TTS. A simple yet effective greedy search algorithm is proposed to determine the optimal configuration. Experimental results show that latency-optimal TTS enables 32B model to achieve 82.3% accuracy on MATH-500 within 1 minute, while 3B model attains notable accuracy within just 10 seconds, showing significant improvements in both speed and accuracy.
@@ 7 Limitations
8 Potential Risks
While our work shows test-time scaling under latency budget and our methods can significantly reduce inference latency, it introduces several risks. For instance, TTS could sometimes degrade performance on underrepresented data domains, exacerbating fairness issues. Scaled models may become more susceptible to adversarial prompts that exploit simplified decision pathways.
9 License For Artifacts
The artifacts utilized in this study, including datasets, codebase, and pre-trained models, are sourced from publicly available repositories under permissive licenses. All datasets adhere to open-access licenses (MIT License), ensuring compliance with redistribution and modification terms. Codebase adheres to open-access licenses (MIT License). For pre-trained models, we verify compatibility with Apache License 2.0. This alignment guarantees ethical reuse while maintaining transparency in our methodology.
10 Information About Use of AI Assistants
In the preparation of this work, we employ AI assistants to assist with refining academic language, and debugging code segments. The AI tools were used solely for improving clarity, grammatical correctness, and syntactic efficiency—tasks analogous to those performed by a human editor or linter. All conceptual contributions, technical claims, and critical analysis remain the authors’ own.
@@ A.2 Additional Results
We present additional experimental results. The results on the influence of branch-wise parallelism is shown in Figure 11. The results of the influence of branch-wise parallelism under different speculative decoding configurations is shown in Figure 12. The results of sequence-wise parallelism under different branch counts are shown in Figure 13..
@@
Models Configuration Accuracy #Req
1 4 16
s1.1-32B $B=1,\gamma=5$ 82.3% 90.5s 91.2s 92.9s
s1.1-32B $B=4,\gamma=5$ 85.6% 91.3s 93.1s 156.8s
s1.1-32B $B=16,\gamma=5$ 83.8% 90.5s 142.7s 576.8s
s1.1-32B $B=64,\gamma=5$ 83.0% 159.8s 602.8s 2621.8s
Llama-3.1-8B-Instruct $B=1,\gamma=4$ 42.3% 7.6s 7.6s 8.2s
Llama-3.1-8B-Instruct $B=4,\gamma=4$ 47.6% 7.6s 8.3s 26.4s
Llama-3.1-8B-Instruct $B=16,\gamma=4$ 63.1% 8.1s 25.8s 95.8s
Llama-3.1-8B-Instruct $B=64,\gamma=4$ 64.9% 24.4s 91.2s 374.5s
Table 4: Latency under different configurations and number of requests (#Req). Memory-bound workloads are marked with gray. As the number of computed tokens per forward grows, the system transits from memory-bound to compute-bound. The latency of the latter can be measured by #tokens.
A.3 Discussion on prefill phase
The whole inference time of LLM is composed of prefilling and decoding. While prefill time is indeed an important overhead of the overall inference, it is not the dominant factor in our case. For example, we have analyzed MATH-500 and find that the question length distribution is highly skewed towards shorter inputs: 88% of questions contain fewer than 360 tokens, and 99% contain fewer than 800 tokens. In contrast, our model generates responses up to 8,192 tokens in length. On modern GPUs, the prefill phase is significantly faster than the autoregressive decoding phase. Given the ratio of input and output lengths in our dataset, the prefill time is negligible of the total latency.
Therefore, while prefill time is technically part of the inference process, its contribution is small. Excluding it does not affect the conclusions.
A.4 Discussion on large batch, multi-requests scenario
We show the experiments as number of branches grows in the table below. The first column: model configuration. The second column: Accuracy. #Req: number of requests. Each cell contains the latency. With more requests, the workload transitions to a compute-bound state. At this point, latency depends on the FLOPs, which is directly proportional to the token count. Consequently, for our latency-aware strategy, once the workload transitions from memory-bound to compute-bound, the trade-off strategies can be directly informed by previous token count strategy.